OpenStudy (lena772):

Arianna kicks a soccer ball off the ground and into the air with an initial velocity of 42 feet per second. Assume the starting height of the ball is 0 feet. Approximately how long does it take until the soccer ball hits the ground again? 1.3 secs 2.6 secs 16.0 secs 27.6 secs I think it's D, but I see a lot of people saying it's B?

ganeshie8 (ganeshie8):

Are you using this equation : H(t) = -16t^2 + vt + s ?

ganeshie8 (ganeshie8):

plug the given info in the equation : initial height = s = 0, v = 42

ganeshie8 (ganeshie8):

H(t) = -16t^2 + 42t

ganeshie8 (ganeshie8):

set it equal to 0 and solve

OpenStudy (lena772):

H(0)=-16*0^2+42*0?

OpenStudy (lena772):

Sorry I disappeared.

OpenStudy (lena772):

@SolomonZelman @thomaster @iPwnBunnies

jimthompson5910 (jim_thompson5910):

The ball will hit the ground when the height, H(t), is equal to 0 So when H(t) = 0 Plug 0 in for H(t) and solve for t H(t) = -16t^2 + 42t 0 = -16t^2 + 42t -16t^2 + 42t = 0 -2t(8t - 21) = 0 I'll let you finish

OpenStudy (jdoe0001):

 \(\bf \text{initial velocity}\\ h = -16t^2+v_ot+h_o \qquad \text{in feet}\\ \\ \quad \\ v_o=\textit{initial velocity of the object}\\ h_o=\textit{initial height of the object}\\ h=\textit{height of the object at "t" seconds} \\ \quad \\ \quad \\ h = -16t^2+42t+0 \implies 0= -16t^2+42t+0\)